OpenStudy (anonymous):

Element A has 15 protons. Determine the number of neutrons in its isotope with mass number 33. A) 15 B) 18 C) 33 D) 48

OpenStudy (abhisar):

Hey there! Welcome to OpenStudy :) Mass number is equal to number of protons + number of neutrons. Let the number of neutrons in the given isotope be equal to A. Then, A+15=33 Now from the above equation find the value of A.

OpenStudy (anonymous):

I Agree its A

OpenStudy (anonymous):

no solve the equation A+15=33 ---> A=33-15=18